OpenStudy (aglna):

tennis ball has a radius of 6.7 centimeters. What is the volume of a tennis ball? Use 3.14 for pi. Round your answer to the nearest hundredth. 157.48 cm3 1259.19 cm3 4188.79 cm3 9634.08 cm3

pooja195 (pooja195):

What shape is a tennis ball? :)

OpenStudy (aglna):

Round a circle

pooja195 (pooja195):

So a sphere :P

OpenStudy (aglna):

Sphere

pooja195 (pooja195):

Volume of a sphere= \[\huge~\rm~\bf~V=\frac{ 4 }{ 3 }(\pi)(r^3)\]

pooja195 (pooja195):

we are given the radius so we can plug it in

OpenStudy (aglna):

So would it be 157.48

pooja195 (pooja195):

Volume of a sphere= \[\huge~\rm~\bf~V=\frac{ 4 }{ 3 }(\pi)(6.7^3)\] ^that will get you the answer Can you show me how you got that?

OpenStudy (aglna):

I tried to do what you said sorry I am new to this type of math

pooja195 (pooja195):

Volume of a sphere= \[\huge~\rm~\bf~V=\frac{ 4 }{ 3 }(3.14)(6.7^3)\] ^that will get you the answer \[\huge~\rm~\bf~6.7^3=6.7*6.7*6.7=? \]

OpenStudy (aglna):

300.763

pooja195 (pooja195):

Volume of a sphere= \[\huge~\rm~\bf~V=\frac{ 4 }{ 3 }(3.14)(6.7^3)\] \[\huge~\rm~\bf~6.7^3=6.7*6.7*6.7=300.763\] \[\huge~\rm~\bf~V=\frac{ 4 }{ 3 }(3.14)(300.763)\] \[\huge~\rm~3.14*300.763=?\]

OpenStudy (aglna):

944.39582

pooja195 (pooja195):

Good Volume of a sphere= \[\huge~\rm~\bf~V=\frac{ 4 }{ 3 }(3.14)(6.7^3)\] \[\huge~\rm~\bf~6.7^3=6.7*6.7*6.7=300.763\] \[\huge~\rm~\bf~V=\frac{ 4 }{ 3 }(3.14)(300.763)\] \[\huge~\rm~3.14*300.763=944.39582\] \[\huge~\rm~\bf~V=\frac{ 4 }{ 3 }(944.39582)\] \[\huge~\rm~\bf~V=\frac{ 4 }{ 3 }*(944.39582)=?\]

OpenStudy (aglna):

1259.19

pooja195 (pooja195):

Yep and that would be our answer :)

OpenStudy (aglna):

Thank you you really helped

pooja195 (pooja195):

haha no problem! ^_^